SPECTROSCOPIC INVESTIGATION OF THE CRYSTAL STRUCTURE OF THE BINARY SYSTEM LiIO_{3}−HIO_{3}

\underline{ALEXANDER BARABASH}, T.GAVRILKO, A.PETROSYAN, G.PUCHKOVSKAYA, AND A.ROSHIN, Institute of Physics, National Academy of Sciences of Ukraine, 46, Prosp. Nauki, 252022 Kiev, Ukraine.

The lattice dynamics of the pseudo-binary system LiIO_{3}−HIO_{3} has been investigated by means of the IR and Raman spectroscopies and the NQR technique. The spectral changes, observed in the 2600 − 400 cm^{−1} region, indicate the formation of variable ”intermolecular” hydrogen bonds O-H⋯O in the system between the quasi molecular halogenate groups IO_{3}^{−}, HIO_{3} and H(IO_{3})_{2}^{−}. By the analysis of low-frequency Raman spectra of the mixed crystals Li_{1−x}H_{x}IO_{3} with varying the concentration x in the range from 0 to 1, the critical x values of HIO_{3} content were evaluated at which the hydrogen bonds of different strength are formed and the different crystal structure of the binary system is observed. Thus for the concentration range 0 − 0.22 the IR absorption spectra show wide and weak O-H stretching bands. That concentration range corresponds to α-modification LiIO_{3} crystal doped by protons (C_{6}^{6} space group). Within the intermediate concentration range 0.06 < x < 0.22 two separate crystals α-LiIO_{3} and 2LiIO_{3}·HIO_{3} (with unknown crystal structure) exist as it follows from the NQR data. For concentrations x > 0.35 the vibrational spectra of the system correspond to Li-doped orthorhombic α-HIO_{3} crystal. From the NQR and Raman data the amorphous (”glass”) phase was detected in the concentration range 0.22 − 0.35. Our results help to solve some contradictions concerning the binary system structure obtained by different experimental methods.